OpenStudy (anonymous):

The midpoint of a line segment with end points as (2, 3) and (4, y2) is (3, 2). What is the value of y2? Please Help

OpenStudy (anonymous):

y2 = 1

OpenStudy (anonymous):

(2+4)/2=3 6/2=3

OpenStudy (anonymous):

(3+y2)/2=2 4=(y2+3) 4-3=y2 1=y2

OpenStudy (anonymous):

thank you!